OpenStudy (mylahb):

Solve the equation by completing the square. x^2+8x=-1

OpenStudy (mylahb):

@KamiBug @cj49

OpenStudy (cj49):

refer to the previous solution and try to solve it.....ill be here to guide u

OpenStudy (mylahb):

x^2+8x+8=-1+8 (x+4)^2=7

OpenStudy (cj49):

half of 8 is 4 then u need to square it i.e 4x4=16

OpenStudy (mylahb):

x^2+8x+4=-1+4 (x+2)^2=3

OpenStudy (cj49):

x^2+8x+16=-1+16

OpenStudy (mylahb):

where does 16 come from?

OpenStudy (cj49):

half of 8 is 4...after taking the half u need to square it and then add it to both the sides

OpenStudy (mylahb):

so is this right.. x^2+8x+16=-1+16 (x+4)^2=15 x+4 pm squared 15

OpenStudy (cj49):

yessssss

OpenStudy (cj49):

it wud be x= -4 pm sqrt 15